OpenStudy (anonymous):

A small publishing company is planning to publish a new book. The production costs will include one-time fixed costs (such as editing) and variable costs (such as printing). There are two production methods it could use. With one method, the one-time fixed costs will total $64,652 , and the variable costs will be $10 per book. With the other method, the one-time fixed costs will total $12,137 , and the variable costs will be $21.25 per book. For how many books produced will the costs from the two methods be the same?

OpenStudy (anonymous):

64,652+10x=12,137+21.25x. I'm not doing it for you :)

OpenStudy (anonymous):

method 1 cost = 64652 + 10x method 2 cost = 12137 + 21.25x equate these and solve for x

OpenStudy (anonymous):

Thank, I can't seem to come up with the answer